# Pathformer: Multi-scale transformers with Adaptive Pathways for Time Series Forecasting

### 04 Feb 2024 v1

@@ Abstract
Transformer-based models have achieved some success in time series forecasting. Existing methods mainly model time series from limited or fixed scales, making it challenging to capture different characteristics spanning various scales. In this paper, we propose multi-scale transformers with adaptive pathways (Pathformer). The proposed Transformer integrates both temporal resolution and temporal distance for multi-scale modeling. Multi-scale division divides the time series into different temporal resolutions using patches of various sizes. Based on the division of each scale, dual attention is performed over these patches to capture global correlations and local details as temporal dependencies. We further enrich the multi-scale transformer with adaptive pathways, which adaptively adjust the multi-scale modeling process based on the varying temporal dynamics in the input time series, improving the prediction accuracy and generalization of Pathformer. Extensive experiments on eleven real-world datasets demonstrate that Pathformer not only achieves state-of-the-art performance by surpassing all current models but also exhibits stronger generalization abilities under various transfer scenarios.
@@ 1 Introduction
Time series forecasting is an essential task for various industries, such as energy, finance, traffic, and cloud computing (Chen et al., 2012; Cirstea et al., 2022b; Qin et al., 2023; Pan et al., 2023). Motivated by its widespread application in sequence modeling and impressive success in various fields such as CV and NLP (Dosovitskiy et al., 2021; Brown et al., 2020), Transformer (Vaswani et al., 2017) receives emerging attention in time series (Wu et al., 2021; Liu et al., 2022c). Despite the growing performance, recent works have started to challenge the existing designs of Transformers for time series forecasting by proposing simpler linear models with better performance (Zeng et al., 2023). While the capabilities of Transformers are still promising in time series forecasting (Nie et al., 2023), it calls for better designs and adaptations to fulfill its potential.
@@
Real-world time series exhibit diverse variations and fluctuations at different temporal scales. For instance, the utilization of CPU, GPU, and memory resources in cloud computing reveals unique temporal patterns spanning daily, monthly, and seasonal scales. This calls for multi-scale modeling (Mozer, 1991; Ferreira et al., 2006) for time series forecasting, which extracts temporal features and dependencies from various scales of temporal intervals. There are two aspects to consider for multiple scales in time series: temporal resolution and temporal distance. Temporal resolution corresponds to how we view the time series in the model and determines the length of each temporal patch or unit considered for modeling. In Figure 1, the same time series can be divided into small patches (blue) or large ones (yellow), leading to fine-grained or coarse-grained temporal characteristics. Temporal distance corresponds to how we explicitly model temporal dependencies and determines the distances between the time steps considered for temporal modeling. In Figure 1, the black arrows model the relations between nearby time steps, forming local details, while the colored arrows model time steps across long ranges, forming global correlations.
@@
To further explore the capability of extracting correlations in Transformers for time series forecasting, in this paper, we focus on the aspect of enhancing multi-scale modeling with the Transformer architecture. Two main challenges limit the effective multi-scale modeling in Transformers. The first challenge is the incompleteness of multi-scale modeling. Viewing the data from different temporal resolutions implicitly influences the scale of the subsequent modeling process (Shabani et al., 2023). However, simply changing temporal resolutions cannot emphasize temporal dependencies in various ranges explicitly and efficiently. On the contrary, considering different temporal distances enables modeling dependencies from different ranges, such as global and local correlations (Li et al., 2019). However, the exact temporal distances of global and local intervals are influenced by the division of data, which is not complete from a single view of temporal resolution. The second challenge is the fixed multi-scale modeling process. Although multi-scale modeling reaches a more complete understanding of time series, different series prefer different scales depending on their specific temporal characteristics and dynamics. For example, comparing the two series in Figure  1, the series above shows rapid fluctuations, which may imply more attention to fine-grained and short-term characteristics. The series below, on the contrary, may need more focus on coarse-grained and long-term modeling. The fixed multi-scale modeling for all data hinders the grasp of critical patterns of each time series, and manually tuning the optimal scales for a dataset or each time series is time-consuming or intractable. Solving these two challenges calls for adaptive multi-scale modeling, which adaptively models the current data from certain multiple scales.
@@ 2 Related Work
Time Series Forecasting. Time series forecasting involves predicting future time series based on historical observations. Statistical modeling methods based on exponential smoothing and its different flavors serve as a reliable workhorse for time series forecasting (Hyndman & Khandakar, 2008; Li et al., 2022a). Among deep learning methods, GNNs are used to model spatial dependency for correlated time series forecasting (Wu et al., 2020; Zhao et al., 2024; Cheng et al., 2024; Miao et al., 2024). RNNs are introduced to model the temporal dependency (Chung et al., 2014; Wen et al., 2017; Cirstea et al., 2019). DeepAR (Rangapuram et al., 2018) uses RNNs and autoregressive methods to predict future short-term series. CNN models use the temporal convolution to extract the sub-series features (Sen et al., 2019; Liu et al., 2022a; Wang et al., 2023). TimesNet (Wu et al., 2023a) transforms the original one-dimensional time series into a two-dimensional space and captures multi-period features through convolution. MLPs (Das et al., 2023; Oreshkin et al., 2020) also show simple and effective applications in this field. Additionally, some methods are incorporating neural architecture search to discover optimal architectures(Wu et al., 2022; 2023b).
@@ 5 Conclusion
In this paper, we propose Multi-Scale Transformers with Adaptive Pathways (Pathformer) for time series forecasting. It integrates multi-scale temporal resolutions and temporal distances by introducing patch division with multiple patch sizes and dual attention on the divided patches, enabling the comprehensive modeling of multi-scale characteristics. Furthermore, adaptive pathways dynamically select and aggregate scale-specific characteristics based on the different temporal dynamics. These innovative mechanisms collectively empower Pathformer to achieve outstanding prediction performance and demonstrate strong generalization capability on several forecasting tasks.

### 20 Feb 2024 v2

# Multi-scale transformers with Adaptive Pathways for Time Series Forecasting

@@ Abstract
Transformer-based models have achieved some success in time series forecasting. Existing methods mainly model time series from limited or fixed scales, making it challenging to capture different characteristics spanning various scales. In this paper, we propose Pathformer, a multi-scale transformer with adaptive pathways. The proposed Pathformer integrates both temporal resolution and temporal distance for multi-scale modeling. Multi-scale division divides the time series into different temporal resolutions using patches of various sizes. Based on the division of each scale, dual attention is performed over these patches to capture global correlations and local details as temporal dependencies. We further enrich the multi-scale transformer with adaptive pathways, which adaptively adjust the multi-scale modeling process based on the varying temporal dynamics in the input time series, improving the prediction accuracy and generalization of Pathformer. Extensive experiments on eleven real-world datasets demonstrate that Pathformer not only achieves state-of-the-art performance by surpassing all current models but also exhibits stronger generalization abilities under various transfer scenarios.
@@ 1 Introduction
Time series forecasting is an essential function for various industries, such as energy, finance, traffic, logistics, and cloud computing (Chen et al., 2012; Cirstea et al., 2022b; Ma et al., 2014; Zhu et al., 2023; Pan et al., 2023; Pedersen et al., 2020), and is also a foundational building block for other time series analytics, e.g., outlier detection Campos et al. (2022); Kieu et al. (2022b). Motivated by its widespread application in sequence modeling and impressive success in various fields such as CV and NLP (Dosovitskiy et al., 2021; Brown et al., 2020), Transformer (Vaswani et al., 2017) receives emerging attention in time series (Wen et al., 2023; Wu et al., 2021; Chen et al., 2022). Despite the growing performance, recent works have started to challenge the existing designs of Transformers for time series forecasting by proposing simpler linear models with better performance (Zeng et al., 2023). While the capabilities of Transformers are still promising in time series forecasting (Nie et al., 2023), it calls for better designs and adaptations to fulfill its potential.
@@
Real-world time series exhibit diverse variations and fluctuations at different temporal scales. For instance, the utilization of CPU, GPU, and memory resources in cloud computing reveals unique temporal patterns spanning daily, monthly, and seasonal scales Pan et al. (2023). This calls for multi-scale modeling (Mozer, 1991; Ferreira et al., 2006) for time series forecasting, which extracts temporal features and dependencies from various scales of temporal intervals. There are two aspects to consider for multiple scales in time series: temporal resolution and temporal distance. Temporal resolution corresponds to how we view the time series in the model and determines the length of each temporal patch or unit considered for modeling. In Figure 1, the same time series can be divided into small patches (blue) or large ones (yellow), leading to fine-grained or coarse-grained temporal characteristics. Temporal distance corresponds to how we explicitly model temporal dependencies and determines the distances between the time steps considered for temporal modeling. In Figure 1, the black arrows model the relations between nearby time steps, forming local details, while the colored arrows model time steps across long ranges, forming global correlations.
@@
To further explore the capability of extracting correlations in Transformers for time series forecasting, in this paper, we focus on the aspect of enhancing multi-scale modeling with the Transformer architecture. Two main challenges limit the effective multi-scale modeling in Transformers. The first challenge is the incompleteness of multi-scale modeling. Viewing the data from different temporal resolutions implicitly influences the scale of the subsequent modeling process (Shabani et al., 2023). However, simply changing temporal resolutions cannot emphasize temporal dependencies in various ranges explicitly and efficiently. On the contrary, considering different temporal distances enables modeling dependencies from different ranges, such as global and local correlations (Li et al., 2019). However, the exact temporal distances of global and local intervals are influenced by the division of data, which is incomplete from a single view of temporal resolution. The second challenge is the fixed multi-scale modeling process. Although multi-scale modeling reaches a more complete understanding of time series, different series prefer different scales depending on their specific temporal characteristics and dynamics. For example, comparing the two series in Figure  1, the series above shows rapid fluctuations, which may imply more attention to fine-grained and short-term characteristics. The series below, on the contrary, may need more focus on coarse-grained and long-term modeling. The fixed multi-scale modeling for all data hinders the grasp of critical patterns of each time series, and manually tuning the optimal scales for a dataset or each time series is time-consuming or intractable. Solving these two challenges calls for adaptive multi-scale modeling, which adaptively models the current data from certain multiple scales.
@@ 2 Related Work
Time Series Forecasting. Time series forecasting predicts future observations based on historical observations. Statistical modeling methods based on exponential smoothing and its different flavors serve as a reliable workhorse for time series forecasting (Hyndman & Khandakar, 2008; Li et al., 2022a). Among deep learning methods, GNNs model spatial dependency for correlated time series forecasting (Jin et al., 2023a; Wu et al., 2020; Zhao et al., 2024; Cheng et al., 2024; Miao et al., 2024; Cirstea et al., 2021). RNNs model the temporal dependency (Chung et al., 2014; Kieu et al., 2022a; Wen et al., 2017; Cirstea et al., 2019). DeepAR (Rangapuram et al., 2018) uses RNNs and autoregressive methods to predict future short-term series. CNN models use the temporal convolution to extract the sub-series features (Sen et al., 2019; Liu et al., 2022a; Wang et al., 2023). TimesNet (Wu et al., 2023a) transforms the original one-dimensional time series into a two-dimensional space and captures multi-period features through convolution. LLM-based methods also show effective performance in this field (Jin et al., 2023b; Zhou et al., 2023). Additionally, some methods are incorporating neural architecture search to discover optimal architectures(Wu et al., 2022; 2023b).
@@ 5 Conclusion
In this paper, we propose Pathformer, a Multi-Scale Transformer with Adaptive Pathways for time series forecasting. It integrates multi-scale temporal resolutions and temporal distances by introducing patch division with multiple patch sizes and dual attention on the divided patches, enabling the comprehensive modeling of multi-scale characteristics. Furthermore, adaptive pathways dynamically select and aggregate scale-specific characteristics based on the different temporal dynamics. These innovative mechanisms collectively empower Pathformer to achieve outstanding prediction performance and demonstrate strong generalization capability on several forecasting tasks.
Acknowledgments
This work was partially supported by Alibaba Innovative Research Program.